# A Unified Evaluation Framework for Multi-Annotator Tendency Learning

### v1

@@ 1. Introduction
In real-world multi-annotation scenarios, such as medical image analysis (Liao et al., 2024), sentiment analysis (Lian et al., 2023), and visual perception (Zhang et al., 2023c), different annotators often provide different labels to the same sample (Raykar et al., 2010) due to varying personal backgrounds, subjective interpretations, and expertise levels. These systematic patterns in labeling behavior—which we term tendencies—represent valuable information about individual perspectives, cognitive biases, and domain expertise.
As illustrated in Figure 1 (top left), the dominant paradigm in multi-annotator learning has been Consensus-oriented Learning (CoL), treating annotator disagreements as noise (Yan et al., 2014; Jinadu et al., 2023) to be averaged away through aggregation techniques like PADL (Liao et al., 2024) and MaDL (Herde et al., 2023). Recent works have emerged in multi-annotator learning that shift focus from CoL to Individual Tendency Learning (ITL) (Figure 1, top right) with sophisticated architectures like QuMAB (Zhang et al., 2025a) and TAX (Cheng et al., 2023) that treat annotator diversity as valuable information to model individual annotators for capturing their unique tendencies and providing behavioral explainability.
@@
Furthermore, existing explainability assessments in multi-annotator systems rely primarily on qualitative analysis, lacking quantitative metrics to evaluate whether learned explanations accurately reflect genuine annotator behavioral relationships. A critical question remains unanswered: How can we systematically evaluate which methods truly preserve individual annotator tendencies and provide meaningful behavioral explanations? To address these fundamental evaluation challenges, we propose a unified evaluation framework that systematically assesses ITL methods across both tendency capture and explainability quality dimensions.
@@ 2.1. Multi-annotator Learning Paradigms
The dominant approach in multi-annotator learning has been consensus-oriented, aggregating diverse annotations into a single ground truth. Methods evolved from simple majority voting to sophisticated techniques using probabilistic modeling (Dawid and Skene, 1979), EM inference (Raykar et al., 2010; Whitehill et al., 2009a), and deep learning (Rodrigues et al., 2014; Albarqouni et al., 2016; Welinder et al., 2010). Advanced approaches model annotator characteristics—reliability, expertise, and confusion patterns—to improve consensus quality (Tanno et al., 2019a; Cao et al., 2019; Chen et al., 2023; Tanno et al., 2019b; Wang et al., 2023). Recent advances explored nuanced annotator modeling: D-LEMA (Mirikharaji et al., 2021) ensembles annotator learners, PADL (Liao et al., 2024) fits Gaussian distributions, and MaDL (Herde et al., 2023) models confusion matrices. SimLabel (Zhang et al., 2025b) addresses the practical challenge of missing labels in multi-annotator scenarios. Despite architectural innovations, these methods remain consensus-oriented, treating disagreements as noise to be averaged away rather than as valuable information to be modeled. The underlying computer vision and machine learning techniques used in multi-annotator learning have broader applications across various domains (Zhang et al., 2023a, b, 2022; Zhang, 2024a; Sha et al., 2024; Zhang et al., 2014), though annotator disagreements in multi-annotator scenarios reflect genuine perspective differences rather than noise. Some works recognize annotator diversity’s value (Snow et al., 2008a), particularly in subjective domains where no absolute ground truth exists. QuMAB (Zhang et al., 2025a) uses learnable queries to model individual patterns with interpretability. However, these efforts lack systematic evaluation frameworks to assess tendency capture and behavioral explanation quality.
@@ 2.2. Tendency Capture Evaluation Metrics
Evaluation in multi-annotator learning has predominantly focused on consensus accuracy, measuring how well methods predict majority votes or expert-defined ground truth (Snow et al., 2008b; Carpenter, 2011). Individual annotator modeling is typically assessed through prediction accuracy or log-likelihood on held-out labels (Raykar et al., 2010; Whitehill et al., 2009b). Traditional agreement metrics—Fleiss’ kappa (Fleiss, 1971a), Krippendorff’s alpha (Gwet, 2011), and Cohen’s kappa (Viera and Garrett, 2005)—measure inter-annotator agreement in raw annotations but do not assess how well computational methods preserve these patterns during learning. Recent approaches focused on annotator modeling quality through likelihood-based metrics (Passonneau and Carpenter, 2014) or behavior correlations (Ding et al., 2023; Gordon et al., 2022). However, these assess individual behaviors rather than inter-annotator relationship structures. Existing metrics either evaluate consensus quality or individual accuracy, but none quantify how well methods maintain inter-annotator relationships that encode perspective diversity. Our proposed Difference of Inter-annotator Consistency (DIC) metric addresses this gap by measuring tendency capture through consistency pattern matching.
@@ 2.3. Multi-annotator Explainability Assessment
Explainability in multi-annotator learning has received limited systematic attention, with most works providing qualitative rather than quantitative evaluation. Existing approaches include feature-based explanations using learned representations (PADL (Liao et al., 2024), MaDL (Herde et al., 2023)) and attention-based explanations showing spatial/temporal patterns (QuMAB (Zhang et al., 2025a), TAX (Cheng et al., 2023)), but rely on subjective interpretation without validation frameworks. Current evaluation faces limitations: broader XAI metrics (Christoph, 2020; Nauta et al., 2023) target single-model scenarios and cannot assess inter-annotator relationships, while human evaluation (Doshi-Velez and Kim, 2017) remains expensive and difficult to standardize. Our proposed Behavior Alignment Explainability (BAE) metric addresses this gap, a quantitative assessment applicable to both explanation types. By comparing explanation-derived similarities with ground-truth behavioral patterns, BAE systematically evaluates whether explanations capture genuine annotator relationships regardless of the underlying mechanism.
@@ 3.1. Difference of Inter-annotator Consistency (DIC)
where normalization ensures scale-invariance across datasets. Lower DIC indicates better tendency capture, with random assignments producing high values (0.86-0.93) due to inconsistent prediction patterns, while effective methods achieve substantially lower scores by maintaining inter-annotator structural relationships.
@@ 3.2. Behavior Alignment Explainability (BAE)
where $\mathbf{A}_{i}^{\text{avg}}=\frac{1}{|\mathcal{S}_{i}|}\sum_{x\in\mathcal{S}_{i}}\text{Attention}(x,i)$ represents the average attention pattern for annotator $i$ across input regions (image patches for STREET, video frames for AMER). This approach provides a complementary perspective by analyzing fine-grained spatial/temporal behavioral patterns, offering different analytical insights rather than consistently superior performance compared to feature-level assessment.
@@ 4. Experiment
Datasets. We evaluate on two multi-annotator datasets: AMER (video-based emotion recognition with 13 annotators) and STREET (urban image impression assessment with 10 annotators across five dimensions: Happiness, Healthiness, Safety, Liveliness, and Orderliness). The AMER dataset’s complexity in capturing temporal emotion dynamics aligns with recent advances in time-sensitive emotion recognition (Zhang, 2024b; Zhang et al., 2024). These datasets provide dense, diverse annotations crucial for modeling annotator-specific behavior patterns.
@@ 4.1. DIC Assessment
We evaluate DIC effectiveness through visual consistency analysis, quantitative comparison across architectures, and validation against standard metrics.
@@
Quantitative Results. Table 1 presents comprehensive DIC scores across four architectures on both datasets. QuMAB consistently achieves the lowest DIC scores, indicating a superior tendency capture across visual and affective domains. Compared to D-LEMA, QuMAB improves DIC by 0.12–0.18 across perspectives, demonstrating robust capture of individualized labeling patterns under varying contexts.
Figure 4. Visualization analysis about behavior alignment explainability (BAE) via 2D projection of annotator representations using Multidimensional Scaling (MDS) on the STREET dataset (safety perspective). Results show progressively improved alignment with higher BAE scores. Region-level QuMAB has a relatively better alignment. Each point denotes an annotator, and proximity indicates higher behavioral similarity. Same colors denote clusters of annotators with strong agreement ($\kappa>0.6$).
@@
Traditional Metrics Comparison. To further highlight the value of DIC, we compare it with three conventional metrics: (1) Accuracy (ACC) for individual annotator prediction quality; (2) Fleiss’ Kappa (FK) (Fleiss, 1971b) measures the overall inter-annotator agreement adjusted for chance; (3) Pearson Correlation Coefficient (PCC) (Benesty et al., 2009) quantifies the similarity between the predicted and true annotation vectors per annotator, reflecting linear trends but not structural agreement across annotators. Table 2 shows results on AMER and STREET (safety perspective). While traditional metrics show limited variation (ranges within 0.06–0.09), DIC reveals more pronounced differences. D-LEMA exhibits moderate traditional scores but the highest DIC, indicating structural modeling failure. QuMAB achieves both the best traditional scores and the lowest DIC, confirming faithful modeling of individual tendencies. These findings confirm that traditional metrics may obscure structural modeling failures, whereas DIC provides clearer, complementary insights by explicitly evaluating tendency capture.
@@ 4.2. BAE Assessment
Behavioral Alignment Analysis. Figure 4 shows MDS projections of learned annotator representations for STREET (safety perspective). Ground truth establishes clear behavioral clusters, with proximity indicating similarity and colors representing high-agreement groups ($\kappa>0.6$). The progression shows: D-LEMA (0.39) with significant distortion, PADL (0.44) with moderate preservation, MaDL (0.47) with better structure, and QuMAB (0.54) with closest ground-truth alignment. QuMAB’s region-level assessment (0.57) complements feature-level analysis by projecting inter-annotator similarities—derived from high-attention input regions—into feature space, offering finer-grained behavioral insight and enhancing overall evaluation robustness. These comparisons highlight each method’s capacity to preserve annotator-specific behavioral relationships through learned representations.
@@ 4.3. Ablation Study
We conduct controlled ablation experiments using carefully designed baseline scenarios representing extreme cases of tendency capture and explainability quality to validate DIC and BAE effectiveness and sensitivity.

### v2

@@ 1. Introduction
In real-world multi-annotation scenarios, such as medical image analysis (Liao et al., 2024), sentiment analysis (Lian et al., 2023), and visual perception (Zhang et al., 2023a), different annotators often provide different labels to the same sample (Raykar et al., 2010) due to varying personal backgrounds, subjective interpretations, and expertise levels. These systematic patterns in labeling behavior—which we term tendencies—represent valuable information about individual perspectives, cognitive biases, and domain expertise.
As illustrated in Figure 1 (top left), the dominant paradigm in multi-annotator learning has been Consensus-oriented Learning (CoL), treating annotator disagreements as noise (Yan et al., 2014; Jinadu et al., 2023) to be averaged away through aggregation techniques like PADL (Liao et al., 2024) and MaDL (Herde et al., 2023). Recent works have emerged in multi-annotator learning that shift focus from CoL to Individual Tendency Learning (ITL) (Figure 1, top right) with sophisticated architectures like QuMAB (Zhang et al., 2025b) and TAX (Cheng et al., 2023) that treat annotator diversity as valuable information to model individual annotators for capturing their unique tendencies and providing behavioral explainability.
@@
Furthermore, existing explainability assessments in multi-annotator systems rely primarily on qualitative analysis, lacking quantitative metrics to evaluate whether learned explanations accurately reflect genuine annotator behavioral relationships. A critical question remains unanswered: How can we systematically evaluate which methods truly preserve individual annotator tendencies and provide meaningful behavioral explanations? To address these fundamental evaluation challenges, we propose a unified evaluation framework that systematically assesses ITL methods across both tendency capture and explainability quality dimensions. Importantly, accurately reproducing individual annotations does not necessarily imply that the relational structure within the annotator pool is preserved. Two models with similar per-annotator accuracy may induce very different inter-annotator similarity structures, leading to divergent interpretations of group dynamics and annotator behaviors. Therefore, evaluating structural consistency among annotators provides complementary information beyond instance-level accuracy.
@@ 2.1. Multi-annotator Learning Paradigms
The dominant approach in multi-annotator learning has been consensus-oriented, aggregating diverse annotations into a single ground truth. Methods evolved from simple majority voting to sophisticated techniques using probabilistic modeling (Dawid and Skene, 1979), EM inference (Raykar et al., 2010; Whitehill et al., 2009a), and deep learning (Rodrigues et al., 2014; Albarqouni et al., 2016; Welinder et al., 2010). Advanced approaches model annotator characteristics—reliability, expertise, and confusion patterns—to improve consensus quality (Tanno et al., 2019b; Cao et al., 2019; Chen et al., 2023; Tanno et al., 2019a; Wang et al., 2023). Recent advances explored nuanced annotator modeling: D-LEMA (Mirikharaji et al., 2021) ensembles annotator learners, PADL (Liao et al., 2024) fits Gaussian distributions, and MaDL (Herde et al., 2023) models confusion matrices. SimLabel (Zhang et al., 2025d, c) addresses the practical challenge of missing labels in multi-annotator scenarios. Despite architectural innovations, these methods remain consensus-oriented, treating disagreements as noise to be averaged away rather than as valuable information to be modeled. The underlying computer vision and machine learning techniques used in multi-annotator learning have broader applications across various domains (Zhang et al., 2023b, c, 2021, 2022; Zhang, 2024a; Sha et al., 2024, 2026; Zhang et al., 2014; Sha et al., 2025; Zhang et al., ), though annotator disagreements in multi-annotator scenarios reflect genuine perspective differences rather than noise. Some works recognize annotator diversity’s value (Snow et al., 2008b), particularly in subjective domains where no absolute ground truth exists. QuMAB (Zhang et al., 2025b, a) uses learnable queries to model individual patterns with interpretability. However, these efforts lack systematic evaluation frameworks to assess tendency capture and behavioral explanation quality.
@@ 2.2. Tendency Capture Evaluation Metrics
Evaluation in multi-annotator learning has predominantly focused on consensus accuracy, measuring how well methods predict majority votes or expert-defined ground truth (Snow et al., 2008a; Carpenter, 2011). Individual annotator modeling is typically assessed through prediction accuracy or log-likelihood on held-out labels (Raykar et al., 2010; Whitehill et al., 2009b). Traditional agreement metrics—Fleiss’ kappa (Fleiss, 1971b), Krippendorff’s alpha (Gwet, 2011), and Cohen’s kappa (Viera and Garrett, 2005)—measure inter-annotator agreement in raw annotations but do not assess how well computational methods preserve these patterns during learning. Recent approaches focused on annotator modeling quality through likelihood-based metrics (Passonneau and Carpenter, 2014) or behavior correlations (Ding et al., 2023; Gordon et al., 2022). However, these assess individual behaviors rather than inter-annotator relationship structures. Existing metrics either evaluate consensus quality or individual accuracy, but none quantify how well methods maintain inter-annotator relationships that encode perspective diversity. Our proposed Difference of Inter-annotator Consistency (DIC) metric addresses this gap by measuring tendency capture through consistency pattern matching.
@@ 2.3. Multi-annotator Explainability Assessment
Explainability in multi-annotator learning has received limited systematic attention, with most works providing qualitative rather than quantitative evaluation. Existing approaches include feature-based explanations using learned representations (PADL (Liao et al., 2024), MaDL (Herde et al., 2023)) and attention-based explanations showing spatial/temporal patterns (QuMAB (Zhang et al., 2025b), TAX (Cheng et al., 2023)), but rely on subjective interpretation without validation frameworks. Current evaluation faces limitations: broader XAI metrics (Christoph, 2020; Nauta et al., 2023) target single-model scenarios and cannot assess inter-annotator relationships, while human evaluation (Doshi-Velez and Kim, 2017) remains expensive and difficult to standardize. Our proposed Behavior Alignment Explainability (BAE) metric addresses this gap, a quantitative assessment applicable to both explanation types. By comparing explanation-derived similarities with ground-truth behavioral patterns, BAE systematically evaluates whether explanations capture genuine annotator relationships regardless of the underlying mechanism.
@@ 3.1. Difference of Inter-annotator Consistency (DIC)
where normalization ensures scale-invariance across datasets. Lower DIC indicates better tendency capture, with random assignments producing high values (0.86-0.93) due to inconsistent prediction patterns, while effective methods achieve substantially lower scores by maintaining inter-annotator structural relationships. While we use Cohen’s for categorical annotations in this work, DIC only requires a well-defined inter-annotator similarity measure and can be readily extended to ordinal or continuous annotations by adopting appropriate agreement or correlation metrics.
@@ 3.2. Behavior Alignment Explainability (BAE)
where $\mathbf{A}_{i}^{\text{avg}}=\frac{1}{|\mathcal{S}_{i}|}\sum_{x\in\mathcal{S}_{i}}\text{Attention}(x,i)$ represents the average attention pattern for annotator $i$ across input regions (image patches for STREET, video frames for AMER). This approach provides a complementary perspective by analyzing fine-grained spatial/temporal behavioral patterns, offering different analytical insights rather than consistently superior performance compared to feature-level assessment. This analysis is architecture-dependent and is not intended as a universally comparable evaluation across all ITL methods.
In practice, feature-level BAE provides a global assessment of tendency alignment and is generally sufficient when annotator behavior differences are primarily semantic. Region-level BAE is particularly informative when annotator tendencies manifest as localized or spatially heterogeneous attention patterns, such as region-specific preferences or biases.
@@ 4. Experiment
Datasets. We evaluate on two representative multi-annotator datasets: AMER (video-based emotion recognition with 13 annotators (Zhang, 2024b; Zhang et al., 2024)) and STREET (urban image impression assessment with 10 annotators across five dimensions: Happiness, Healthiness, Safety, Liveliness, and Orderliness), which cover complementary settings in terms of modality (video vs. image) and annotation semantics, and provide dense, diverse annotations crucial for modeling annotator-specific behavior patterns. Our focus is on evaluating Individual Tendency Learning (ITL) methods under fair and controlled conditions; thus, experiments are restricted to domains where multiple representative ITL baselines are available. Note that the proposed framework itself is modality-agnostic and can be directly extended to other domains when suitable ITL models are available.
@@ 4.1. DIC Assessment
We evaluate DIC effectiveness through visual consistency analysis, quantitative comparison across architectures, and validation against standard metrics. It is important to note that DIC evaluates the structural alignment of inter-annotator relationships, which is complementary to instance-level accuracy metrics such as ACC. A model may achieve similar DIC values while differing in ACC, reflecting different trade-offs between tendency structure modeling and pointwise prediction accuracy. Therefore, DIC is not intended to replace traditional accuracy metrics, but to be jointly considered with them for a more comprehensive evaluation.
@@
Quantitative Results. Table 1 presents comprehensive DIC scores across four architectures on both datasets. QuMAB consistently achieves the lowest DIC scores, indicating a superior tendency to capture across visual and affective domains. Compared to D-LEMA, QuMAB improves DIC by 0.12–0.18 across perspectives, demonstrating robust capture of individualized labeling patterns under varying contexts.
Figure 4. Visualization analysis about behavior alignment explainability (BAE) via 2D projection of annotator representations using Multidimensional Scaling (MDS) on the STREET dataset (safety perspective). Results show progressively improved alignment with higher BAE scores. QuMAB-region provides a complementary view by incorporating attention-over-region patterns. Each point denotes an annotator, and proximity indicates higher behavioral similarity. Same colors denote clusters of annotators with strong agreement ($\kappa>0.6$).
@@
Traditional Metrics Comparison. To further highlight the value of DIC, we compare it with three conventional metrics: (1) Accuracy (ACC) for individual annotator prediction quality; (2) Fleiss’ Kappa (FK) (Fleiss, 1971a) measures the overall inter-annotator agreement adjusted for chance; (3) Pearson Correlation Coefficient (PCC) (Benesty et al., 2009) quantifies the similarity between the predicted and true annotation vectors per annotator, reflecting linear trends but not structural agreement across annotators. Table 2 shows results on AMER and STREET (safety perspective). While traditional metrics show limited variation (ranges within 0.06–0.09), DIC reveals more pronounced differences. D-LEMA exhibits moderate traditional scores but the highest DIC, indicating structural modeling failure. QuMAB achieves both the best traditional scores and the lowest DIC, confirming faithful modeling of individual tendencies. These findings confirm that traditional metrics may obscure structural modeling failures, whereas DIC provides clearer, complementary insights by explicitly evaluating tendency capture.
@@ 4.2. BAE Assessment
Behavioral Alignment Analysis. Figure 4 shows MDS projections of learned annotator representations for STREET (safety perspective). Ground truth establishes clear behavioral clusters, with proximity indicating similarity and colors representing high-agreement groups ($\kappa>0.6$). The progression shows: D-LEMA (0.39) with significant distortion, PADL (0.44) with moderate preservation, MaDL (0.47) with better structure, and QuMAB (0.54) with closest ground-truth alignment. QuMAB’s region-level assessment (0.57) complements feature-level analysis by projecting inter-annotator similarities—derived from high-attention input regions—into feature space, offering finer-grained behavioral insight and providing complementary fine-grained behavioral insights. These comparisons highlight each method’s capacity to preserve annotator-specific behavioral relationships through learned representations.
@@ 4.3. Ablation Study
We conduct controlled ablation experiments using carefully designed baseline scenarios representing extreme cases of tendency capture and explainability quality to validate DIC and BAE effectiveness and sensitivity. These simulated scenarios are designed as controlled reference points to illustrate the sensitivity and discriminative behavior of the proposed metrics, rather than to serve as exhaustive experimental benchmarks.